# VILA2: VILA Augmented VILA

### v1

@@ 3 Experiments
Model Architecture. We follow the architecture from VILA [25], where a multi-modal large model consists of three key components: an LLM for auto-regressive generation, a visual encoder for extracting visual features, and an image-text projector to align the visual and text modalities.
We use Llama2-7B [9] for exploratory experiments to address the question “To what extent can a VLM self-bootstrap?”. Then we switch to our previous SOTA settings with Llama3-8B-Instruct [33] and Yi-34B [14] when compared to other methods. For visual encoders, we use SigLIP [34] for LLaMA-series models and InternViT-6B [35] for the Yi-34B model. For projection layers, we follow LLaVA [16, 36] to adapt simple linear layers for bridging image and text modalities. At the same time, we introduce a 4$\times$ downsampling of visual tokens by concatenating $2\times 2$ neighboring patches along the channel dimension and using a simple MLP for the downsampling process.
@@
    Projector Initialization. The language models and ViT are separately pre-trained, while the projector is randomly initialized. To initially align the feature space between the visual and text modalities, we utilize the LLaVA align dataset [16].
@@
    Vision-Language Pre-training. We then pre-train the model (LLM and the projector) on the visual language corpus. We consider interleaved image text corpus (e.g., MMC4 [26]) and image-text pairs (e.g., COYO [28]). We apply our VILA2 for the pre-training data and the augmented data would used in this stage to replace original COYO captions.
@@
    Visual Instruction-tuning. Finally, we perform instruction tuning of the pre-trained model on visual language instruction datasets. The blending details is attached in the appendix.
Without specifically mentioned, our experiments are conducted with 128 GPUs and a global batch size of 1024. We employ AdamW optimizer with learning rate {$10^{-3}$, $5\times 10^{-5}$, $2\times 10^{-5}$} for aforementioned three stages respectively. Each stage is trained with one epoch with a 3% warmup strategy. No weight decay is applied. In some self-/specialist augmented training, VILA2 may involve extra stage to further improve. Please refer to Section. 2.2 and Appendix for more details.
Data. Our pretraining stage consists of 6M sampled MMC4 [26], 25M sampled Coyo [28], and the full ShareGPT4v [30]. To ensure a fair comparison, we only replace the text captions during our experiments while keeping all image sources the same. We use two SFT data blends for different purposes: a smaller blend of 1.8M samples for exploratory experiments in Table 3–Table 4; a larger blend of 5.9M samples augmented from VILA’s training receipt***https://github.com/NVlabs/VILA/tree/main/data_prepare, for state-of-the-art experiments in Table 5–Table 6. Detailed SFT recipe can be found in the appendix.
Specialist Setups.
  1. 1.
    Spatial Specialists. We continued training the specialist from the stage 2, ALLaVA caption [37], and GPT-4V caption from ShareGPT4V [30].
  2. 2.
    Grounding Specialist. We split the cleaned 4M grounded narration into Grounding-Short 3M and Grounding-Long 838K for a two-stage training process. In stage 3, we combined Grounding-Short 3M with ALLaVA caption [37] to adapt to new tasks of grounded narration while maintaining the narrative ability. In stage 4, we combine Grounding-Long 838K with Shikra GPT-4 [38], Visual7W [39], LLaVA-SFT, and 100K GPT-4V captions from ShareGPT4V to sustain both narrative and instruction following capacities.
  3. 3.
    OCR Specialist. We trained our OCR specialist with various internet datasets focused on text recognition, understanding, and reasoning, including LLaVA-SFT, TextOCR-GPT4V [40], SynthDoG-En [41], OCRVQA [42], TextCaps [43], ArxivQA [44], DocVQA [45], AI2D [46], ChartQA [47], LLaVAR [48] and 35 OCR-related datasets from The Cauldron [49].
@@ 3.1 Self-Augmentation Results
Figure 4: VILA2 continuously enhances caption quality over self-augmenting. The sample is from the COYO [28]. We mark facts in green, hallucinations in red, and spatial information in blue.
VILA2 Enriches Dataset Text Quality. The main VLM’s performance boost comes from the caption quality improvement. As demonstrated in Table. 1, the average number of words grows quickly after self-augmenting and saturates around round-3 and round-4. This is consistent with the trend observed in the benchmark side. In round-1, caption length increases significantly from 12 to 101, while in round-3, the increase is only from 117 to 126. The increase in caption length is no longer significant after round-1 but we still observe consistent improvement on VLM benchmarks. We hypothesize that self-augmentation after round-1 primarily improves caption quality, providing more accurate details and reducing hallucination. The results are visualized in Figure. 4. The origin caption for the provided example is brief and short, only describing Boris’ riding action and no other information. After self-augmenting once, the caption becomes longer and much more informative, including descriptions of dressing and surrounding environments. Several new hallucinations are also added to the text (e.g., suit color, basket not showing in the image). While self-augmenting iterates, the hallucinations gradually get removed from captions. For example, the no-showing basket and wrongly-read "Barcardi" text are no longer mentioned in round-3. $\text{VILA}_{3}$ only describe visual elements that are confident with.
VILA2 Bootstraps VILA’s Performance. We follow the same pre-training + SFT process as VILA [25] and sample 5% data from the pre-training phase to ablate. The images are from the existing COYO [28] and MMC4 [26] and in each loop, we use the models trained last round to generate new captions for half of the sampled COYO images. MMC4 is not re-captioned because of its interleaved feature. Other settings are kept the same. We compare $\text{VILA}_{i}$ from different looping steps on common VLM benchmarks. We notice that self-augmented data help improves the model performance across different iterations: $\text{VILA}_{i+1}$ is consistently better than $\text{VILA}_{i}$ and the looping progressively boosts the performance ($\text{VILA}_{1-4}$ in Table. 3). The self-augmenting technique is consistently useful until three rounds. $\text{VILA}_{4}$ reaches saturation and no longer bring consistent improvement of $\text{VILA}_{3}$.
@@
Surpassing the Limit with VILA2 Specialist. The “self-augmentation then training” cycle is yielding a diminishing return and is likely to reach saturation within three iterations, as illustrated in Tab. 3. However, by incorporating external spatial relations data and fine-tuning a spatial specialist VLM, we can extend this process beyond its current limits. Training a specialist VLM with domain-specific knowledge and using task-specific prompts during the re-captioning process can effectively enrich the resulting captions with expertise knowledge, thereby enhancing benchmark performance.
We show that $\text{VILA}_{3}$ spatial relations understanding specialist represents an extra focus on describing the spatial information in Fig. 4. Caption augmented by the specialist (the last example) will keep the most visible details, and provide more information about spatial relations compared to the “Round-4” caption. This additional information encompasses not only the object-to-object relations but also the localization and clear pose of the major focus, which is not present in the SpatialRelationQA dataset. We hypothesize that this improvement might result from the compositional combination of knowledge in other parts of the training data, as a synergy result from both the new task learning and the self-bootstrapping procedure. The effectiveness of these enriched captions is demonstrated in Table  3. Here, we integrate an additional 5% of COYO data re-captioned by the $\text{VILA}_{3}$ specialist into the pre-training stage. Following the same SFT stage, we observe notable improvements in 5 out of the 8 benchmarks.
3.2 Specialist Augmentation Results
We next transition to our previous state-of-the-art settings to explore the significance of scaling-up our VILA2 approach with more pretraining data and larger models to fully study the benefits of specialists. Following our VILA-1.5 release practicehttps://github.com/NVlabs/VILA, we use S2 [58] with Llama 3-8B-Instruct to enhance the visual encoder by scaling on scales rather than increasing input resolution. Additionally, we utilize a larger visual encoder, the Intern ViT-6B, which has a comparable number of parameters to the Yi-34B model to ensure balanced training.
We demonstrate the effectiveness of each specialist on a $\sim$10subset of the entire pretraining data blend and the 1.8M SFT blend, as shown in Table 4. Generally, specialist VLMs show overall improvements across most VQA benchmarks. We then combine annotations from all three specialists into multi-round QA pairs for each image and re-train VILA. This synergy among the specialists proves highly effective, as scaling-up to the full pretraining set yields significant improvements.
@@ 3.2 Specialist Augmentation Results
BLIP-2 [20] Vicuna-13B 224 41.0 41 19.6 61 42.5 46.4 38.1 22.4
InstructBLIP [59] Vicuna-7B 224 49.2 34.5 60.5 50.1 36 23.7 53.4 60.9 26.2
InstructBLIP [59] Vicuna-13B 224 49.5 33.4 63.1 50.7 58.2 25.6
Qwen-VL [22] Qwen-7B 448 78.8 59.3 35.2 67.1 63.8 38.2 7.4 56.3
Qwen-VL-Chat [22] Qwen-7B 448 78.2 57.5 38.9 68.2 61.5 60.6 56.7 58.2
LLaVA-1.5 [60] Vicuna-1.5-7B 336 78.5 62.0 50.0 66.8 58.2 64.3 58.3 58.6 63.4 30.5
LLaVA-1.5 [60] Vicuna-1.5-13B 336 80.0 63.3 53.6 71.6 61.3 67.7 63.6 61.6 70.7 35.4
VILA-7B [25] Llama 2-7B 336 79.9 62.3 57.8 68.2 64.4 68.9 61.7 61.1 69.7 34.9
VILA-13B [25] Llama 2-13B 336 80.8 63.3 60.6 73.7 66.6 70.3 64.3 62.8 73.0 38.8
LLaVA-NeXT-8B [36] Llama 3-8B 672 65.2 72.8 64.6 72.1 80.1
Cambrian-1-8B [61] Llama 3-8B 1024 64.6 80.4 71.7 75.9
Mini-Gemini-HD-8B [62] Llama 3-8B 1536 64.5 75.1 70.2 72.7
@@
Table 5: Comparison with state-of-the-art methods on 10 visual-language benchmarks. Our models consistently improve VILA under a head-to-head comparison, using the same prompts and the same base LLM, showing the effectiveness of enhanced pretraining data quality. We mark the best performance bold and the second-best underlined.
@@ 3.3 Benchmark Comparison to Prior Work
Method Overall Art & Design Business Science Health Human & Social Tech. & Eng.
GPT-4V [23] 55.7 65.3 64.3 48.4 63.5 76.3 41.7
SenseChat-V [63] 50.3 62.7 44.1 42.3 55.7 74.7 43.5
VILA2-40B (ours) 47.9 62.0 42.3 38.5 51.9 71.9 42.3
Qwen-VL-MAX [15] 46.8 64.2 39.8 36.3 52.5 70.4 40.7
InternVL-Chat-V1.2 [35] 46.2 62.5 37.6 37.9 49.7 70.1 40.8
LLaVA-1.6 [36] 44.7 58.6 39.9 36.0 51.2 70.2 36.3
Marco-VL-Plus* 44.3 57.4 34.7 38.5 48.7 72.2 36.7
Yi-VL [14] 41.6 56.1 33.3 32.9 45.9 66.5 36.0
Qwen-VL-PLUS [15] 40.8 59.9 34.5 32.8 43.7 65.5 32.9
Marco-VL-Plus* 40.4 56.5 31.0 31.0 46.9 66.5 33.8
Weito-VL-1.0* 38.4 56.6 30.5 31.1 38.4 59.0 34.2
VILA2-8B (ours) 38.3 54.3 32.0 29.3 39.7 56.8 34.4
InfiMM-Zephyr [64] 35.5 50.0 29.6 28.2 37.5 54.6 31.1
SVIT [65] 34.1 48.9 28.0 26.8 35.5 50.9 28.8
Emu2-Chat [66] 34.1 50.6 27.7 28.0 32.4 50.3 31.3
BLIP-2 FLAN-T5-XXL [20] 34.0 49.2 28.6 27.3 33.7 51.5 30.4
InstructBLIP-T5-XXL [59] 33.8 48.5 30.6 27.6 33.6 49.8 29.4
LLaVA-1.5 [16] 33.6 49.8 28.2 25.9 34.9 54.7 28.3
Table 6: Comparison with state-of-the-art methods on the MMMU dataset. *: model on leader-board with unidentified reference. The models are ranked by overall scores, including both proprietary and open-sourced models. We highlight our results with color green. VILA2 achieves SOTA performance in the open source category.
@@
We now perform a comprehensive comparison to prior work over 10 major benchmarks and summarize results in Table 5 and Table 6. Note that we used a total of 25 million COYO data sampled from the 700 million with the highest CLIP score. We augment the original short real labels with multi-round QA pairs annotated by three specialists, all from 8B models. For 40B models, we continue to train from the stage 2 checkpoints with a mix of 3.75 M recaptioned COYO and a 200K caption dataset [37]. We observed the improvements in quality is consistent and can scale to 40B VILA checkpoints. The detailed training recipes of our 8B and 40B checkpoints are included in the Appendix and will be released jointly with the code base.
@@
Remarkably we observed the enhanced self-augmentation and specialist augmentation training recipes, backed by enhanced and refined datasets, support VILA2 to further push the performance boundary of VILA [25] by noticeable margins across almost all benchmarks, consistent with the ablated performance boosts we observed in previous analysis of Table 3. Moreover, VILA2 now constitutes a SOTA performance on the main MMMU [3] test dataset leaderboard across all open-sourced models, without the usage of proprietary datasets and only based on publicly available datasets.
4 Related Work
Large language models (LLMs). The LLMs arguably began with the introduction of transformers [67] and subsequent efforts to scale them. For the decoder-only paradigm, OpenAI introduced the Generative Pre-trained Transformer (GPT) models [68][6], from GPT-2 (1.5B parameters) to GPT-4 [23] (1.76T), demonstrating that scaling parameters along with high-quality data can produce coherent and contextually relevant text across various domains. Transformer-XL [5] extended the context window, enabling better comprehension of longer texts.
Open-sourced LLMs significantly advanced the field including LLaMa-1/2/3 [8, 9, 69], its instruct-finetuned extension Alpaca [10], and its chatbot version Vicuna [11]. Various models have further pushed the boundaries of LLM capabilities, such as Mistral [12] for generating coherent long-form text, Falcon [70] scaling open-sourced LLMs to 180B parameters, Yi [14] and DeepSeek [71] with applications in supervised fine-tuning for unlocking multilingual capabilities, DBRX [72] with a fine-grained mixture-of-experts architecture, and Gemini [2] with 10M context window. The scale of models is also rapidly increasing, e.g., the LLaMa-3 [69] at 400B parameters. Our work utilizes LLaMa-2-7B [9], LLaMa-3-8B [69] and Yi-34B [14] as the LLM backbones.
Visual language models (VLM). Visual language models have rapidly progressed in recent years. The success mainly comes from pre-training visual and language models on the internet-scale data. CLIP [73] trained a shared feature space of vision and language modalities through contrastive learning on image-caption data. BLIP [74] and BLIP-2 [20] improved CLIP by leveraging noisy pseudo-labels and expanded multi-modal capabilities by aligning the pre-trained encoder to large language models. LLaVA [16] and InstructBLIP [59] demonstrated that jointly training from a collection of diverse datasets as an instruction-following task leads to strong generalization across tasks. Kosmos-2 [32] and PaLI-X [19] largely scaled the pre-training data by pseudo-labeling bounding boxes from performant open-vocabulary object detectors (GLIP [75] and OWL-v2 [76], respectively). They examined that strong perception capabilities such as object detection and OCR translate to better high-level reasoning tasks like visual question-answering (VQA). Our work expands the horizon of data-scaling through our self-augmenting paradigm. Instead of using VLMs directly for data synthesizing, specialist models with domain expertise such as object detection, spatial relation, OCR, and more have a large potential for participating in the data collection pipeline. In this paper, We explore the potential of labeling the internet-scale image data with both self-augmentation techniques, and also a model specialized in spatial relations.

### v2

@@ 3 Experiments
Model Architecture. We follow the architecture from VILA (Lin et al., 2024), where a multi-modal large model consists of three key components: an LLM for auto-regressive generation, a visual encoder for extracting visual features, and an image-text projector to align the modalities.
We use Llama2-7B (Touvron et al., 2023b) for exploratory experiments to address the question “To what extent can a VLM self-bootstrap?”. Then we switch to our previous SOTA settings with Llama3-8B-Instruct (Author, s) and Yi-34B (yi, 2023) when compared to other methods. For visual encoders, we use SigLIP (Zhai et al., 2023) for LLaMA-series models and InternViT-6B (Chen et al., 2023d) for the Yi-34B model. For projection layers, we follow LLaVA (Liu et al., 2023c; 2024c) to adapt simple linear layers for bridging image and text modalities. At the same time, we introduce a 4$\times$ downsampling of visual tokens by concatenating $2\times 2$ neighboring patches along the channel dimension and using a simple MLP for the downsampling process.
Figure 4: VILA2 continuously enhances caption quality over self-augmenting. The sample is from the COYO. We mark facts in green, hallucinations in red, and spatial information in blue. (Please zoom out for the best viewing experience)
@@
    Projector Initialization. The language models and ViT are separately pretrained, while the projector is randomly initialized. To initially align the feature space between the visual and text modalities, we utilize the LLaVA align dataset (Liu et al., 2023c).
@@
    Vision-Language Pretraining. We then pretrain the model (LLM and the projector) on the visual language corpus. We consider interleaved image text corpus (e.g., MMC4 (Zhu et al., 2023b)) and image-text pairs (e.g., COYO (Byeon et al., 2022)). We apply our VILA2 for the pretraining data and the augmented data will be applied in this stage to replace original COYO captions.
@@
    Visual Instruction-tuning. Finally, we perform instruction tuning of the pretrained model on visual language instruction datasets. The blending details is attached in the appendix.
Without specifically mentioned, our experiments are conducted with 128 GPUs and a global batch size of 1024. We employ AdamW optimizer with learning rate {$10^{-3}$, $5\times 10^{-5}$, $2\times 10^{-5}$} for aforementioned three stages respectively. Each stage is trained with one epoch with a 3% warmup strategy. No weight decay is applied. In some self-/specialist augmented training, VILA2 may involve extra stage to further improve. Please refer to Section. 2.2 and Appendix A.5 for more details.
Data. Our pretraining stage consists of 6M sampled MMC4 (Zhu et al., 2023b), 25M sampled Coyo (Byeon et al., 2022), and the full ShareGPT4V (Chen et al., 2023b). To ensure a fair comparison, we only replace the text captions during our experiments while keeping all image sources the same. We use two SFT data blends for different purposes: a smaller blend of 1.8M samples for exploratory experiments in Table 3–Table 4; a larger blend of 5.9M samples augmented from VILA’s training receipt, for SOTA experiments in Table 5–Table 6. Detailed SFT recipe and specialist data fulllist can be found in the Appendix. A.3-Appendix. A.4.
@@ 3.1 Self-Augmentation Results
VILA2 Enriches Dataset Text Quality. The main VLM’s performance boost stems from improved caption quality. As shown in Table 1, caption length increases rapidly after self-augmentation and plateaus around rounds 3 and 4. This aligns with the trend observed in the benchmark results (round 1: 12 to 101, round 3: 117 to 126). Though Caption length does not increase significantly after round-1, we continue to observe consistent improvement on VLM benchmarks. We hypothesize that self-augmentation beyond round-1 primarily enhances caption quality by providing more accurate details and reducing hallucination, as visualized in Figure 4. The initial brief caption is brief and short (only describing Boris’s riding action). The later captions evolves to include more details about clothing and surroundings. Although early iterations may contain some hallucinations (such as a non-existent basket and misread "Barcardi" text), subsequent iterations refine the caption to include only visual elements that can be confidently identified (more evidence in Table 3.4-Table 3.4).
VILA2 Bootstraps VLM’s Performance. We follow the same pretraining + SFT process as VILA (Lin et al., 2024) and sample 5% data from the pretraining phase to ablate. The images are from the existing COYO (Byeon et al., 2022) and MMC4 (Zhu et al., 2023b) and in each loop, we use the models trained last round to generate new captions for half of the sampled COYO images. MMC4 is not re-captioned because of its interleaved feature. Other settings are kept the same. We compare $\text{VILA}_{i}$ from different looping steps on common VLM benchmarks. We notice that self-augmented data help improves the model performance across different iterations: $\text{VILA}_{i+1}$ is consistently better than $\text{VILA}_{i}$ and the looping progressively boosts the performance ($\text{VILA}_{1-4}$ in Table 3). The self-augmenting technique is consistently useful until three rounds. $\text{VILA}_{4}$ reaches saturation and no longer bring consistent improvement of $\text{VILA}_{3}$.
@@
BLIP-2 (Li et al., 2023b) Vicuna-13B 224 41.0 41 19.6 61 42.5 46.4 38.1 22.4
InstructBLIP (Dai et al., 2023) Vicuna-7B 224 49.2 34.5 60.5 50.1 36 23.7 53.4 60.9 26.2
InstructBLIP (Dai et al., 2023) Vicuna-13B 224 49.5 33.4 63.1 50.7 58.2 25.6
Qwen-VL (Bai et al., 2023b) Qwen-7B 448 78.8 59.3 35.2 67.1 63.8 38.2 7.4 56.3
Qwen-VL-Chat (Bai et al., 2023b) Qwen-7B 448 78.2 57.5 38.9 68.2 61.5 60.6 56.7 58.2
LLaVA-1.5 (Liu et al., 2023b) Vicuna-1.5-7B 336 78.5 62.0 50.0 66.8 58.2 64.3 58.3 58.6 63.4 30.5
LLaVA-1.5 (Liu et al., 2023b) Vicuna-1.5-13B 336 80.0 63.3 53.6 71.6 61.3 67.7 63.6 61.6 70.7 35.4
VILA-7B (Lin et al., 2024) Llama 2-7B 336 79.9 62.3 57.8 68.2 64.4 68.9 61.7 61.1 69.7 34.9
VILA-13B (Lin et al., 2024) Llama 2-13B 336 80.8 63.3 60.6 73.7 66.6 70.3 64.3 62.8 73.0 38.8
LLaVA-NeXT-8B (Liu et al., 2024c) Llama 3-8B 672 65.2 72.8 64.6 72.1 80.1
Cambrian-1-8B (Tong et al., 2024) Llama 3-8B 1024 64.6 80.4 71.7 75.9
Mini-Gemini-HD-8B (Li et al., 2024b) Llama 3-8B 1536 64.5 75.1 70.2 72.7
@@
Table 5: Comparison with state-of-the-art methods on 10 visual-language benchmarks. Our models consistently improve VILA under a head-to-head comparison, showing the effectiveness of enhanced pretraining data quality. We mark the best performance bold and the second-best underlined.
3.2 Specialist Augmentation Results
The “self-augmentation then training” cycle reaches a plateau after three iterations, as illustrated in Table 3. However, by incorporating tasks-specific specialists, we can overcome the limit and introduce further performance improvement.
Surpassing the Limit with VILA2 Specialist. The caption augmented by the specialist (the last example) retains the most visible details and provides more information about spatial relations compared to the "Round-4" caption (Figure 4). This additional information includes object-to-object relations, as well as localization and clear pose of the major focus, which are not present in the SpatialRelationQA dataset. We hypothesize this improvement might result from combining knowledge in specialist data and VLM’s training data. The effectiveness of these enriched captions is demonstrated in Table 3 ($\text{VILA}_{3}$ + spatial specialist). Following the same SFT stage, we observe notable improvements in 5 out of the 8 benchmarks.
The More Specialists, The Better The Performance. We next explore the significance of scaling up our VILA2 specialists with more pretraining data using recent state-of-the-art settings. We demonstrate the effectiveness of each specialist using S2 (Shi et al., 2024) with Llama 3-8B-Instruct (meta, 2024). On a 10% subset of pretraining data, specialist VLMs show overall improvements across most VQA benchmarks in Table 4’s first part. We then combine annotations from all three specialists into multi-round QA pairs for each image and retrain VILA. This synergy among the specialists proves highly effective, with scaling up to the full pretraining set yielding significant improvements. Results on larger models will be discussed in next section.
@@ 3.3 Benchmark Comparison to Prior Work
We now perform a comprehensive comparison to prior work over 10 major benchmarks and summarize results in Table 5 and Table 6. Note that we used a total of 25 million COYO data sampled from the 700 million with the highest CLIP score. We augment the original short real labels with multi-round QA pairs annotated by three specialists, all from 8B models. For 40B models, we continue to train from the stage 2 checkpoints with a mix of 3.75 M recaptioned COYO and a 200K caption dataset (Chen et al., 2024). We observed the improvements in quality is consistent and can scale to 40B VILA checkpoints. The detailed training recipes of our 8B and 40B checkpoints are included in the Appendix and will be released jointly with the code base.
@@
Remarkably we observed the enhanced self-augmentation and specialist augmentation training recipes, backed by enhanced and refined datasets, support VILA2 to further push the performance boundary of VILA (Lin et al., 2024) by noticeable margins across almost all benchmarks, consistent with the ablated performance boosts we observed in previous analysis of Table 3. Moreover, VILA2 now constitutes a SOTA performance on the main MMMU (Yue et al., 2024) test dataset leaderboard across all open-sourced models, without the usage of proprietary datasets and only based on publicly available datasets.
Method Overall (Test/Val) Art & Design Business Science Health Human & Social Tech. & Eng.
GPT-4V (OpenAI, 2023a) 56.1 / 56.8 65.3 64.3 48.4 63.5 76.3 41.7
SenseChat-V (Sensetime, 2024) 50.3 / 54.6 62.7 44.1 42.3 55.7 74.7 43.5
VILA2-40B (ours) 47.9 / 53.0 62.0 42.3 38.5 51.9 71.9 42.3
Qwen-VL-MAX (Bai et al., 2023a) 46.8 / 51.4 64.2 39.8 36.3 52.5 70.4 40.7
InternVL-Chat-V1.2 (Chen et al., 2023d) 46.2 / 51.6 62.5 37.6 37.9 49.7 70.1 40.8
Cambrian-1-34B (Liu et al., 2024c) - / 49.7 - - - - - -
LLaVA-1.6 (Liu et al., 2024c) 44.7 / 48.1 58.6 39.9 36.0 51.2 70.2 36.3
Mini-Gemini-HD-34B (Liu et al., 2024c) - / 48.0 - - - - - -
Marco-VL-Plus* 44.3 / 46.2 57.4 34.7 38.5 48.7 72.2 36.7
Yi-VL (yi, 2023) 41.6 / 45.9 56.1 33.3 32.9 45.9 66.5 36.0
Qwen-VL-PLUS (Bai et al., 2023a) 40.8 / 45.2 59.9 34.5 32.8 43.7 65.5 32.9
Marco-VL-Plus* 40.4 / 41.2 56.5 31.0 31.0 46.9 66.5 33.8
Weitu-VL-1.0* 38.4 / - 56.6 30.5 31.1 38.4 59.0 34.2
VILA2-8B (ours) 38.3 / 40.8 54.3 32.0 29.3 39.7 56.8 34.4
InfiMM-Zephyr (Team, 2024) 35.5 / 39.4 50.0 29.6 28.2 37.5 54.6 31.1
SVIT (Zhao et al., 2023) 34.1 / 38.0 48.9 28.0 26.8 35.5 50.9 28.8
Emu2-Chat (Sun et al., 2023) 34.1 / 36.3 50.6 27.7 28.0 32.4 50.3 31.3
BLIP-2 FLAN-T5-XXL (Li et al., 2023b) 34.0 / 35.4 49.2 28.6 27.3 33.7 51.5 30.4
InstructBLIP-T5-XXL (Dai et al., 2023) 33.8 / 35.7 48.5 30.6 27.6 33.6 49.8 29.4
LLaVA-1.5 (Liu et al., 2023c) 33.6 / 36.4 49.8 28.2 25.9 34.9 54.7 28.3
Table 6: Comparison with state-of-the-art methods on the MMMU dataset. *: model on leaderboard with unidentified reference. The models are ranked by overall test set scores (we report scores in a test/validation manner), including both proprietary and open-sourced models. We highlight our results with color green. VILA2 achieves SOTA performance in the open source category.
3.4 Gauging on Synthetic Data Quality and Hallucination
Figure 2 presents cross-evaluation results with Gemini and GPT-4V, showing their increased preference for captions generated in later rounds of self-augmentation. We also provide evidence from an out-of-distribution benchmark and a human blend quality ranking, demonstrating that hallucinations do not increase in these later rounds.
Left-out Benchmark Results. We first select the Visual Spatial Reasoning (VSR (Liu et al., 2023a)) benchmark which does not appear in our training set. This benchmark consists of triplets containing an image, a spatial-focused expression, (e.g., the cow is ahead of the person), and a True or False label indicating its correctness. We observed reduced hallucinations on the VSR benchmark with more iterations of self-augmentation and have not yet reached a plateau, results are shown in Table 3.4.
Human Judge Ranking Test. We further add a more rigorous human test that compares the (re-)captions of 200 randomly sampled images from 11 human evaluators (most are PhD students). These evaluators were tasked with determining which caption exhibits fewer hallucinations, without any knowledge of the sources of the captions. We calculated the win rates of the captions from later self-augmentation rounds against those from earlier rounds, as detailed in Table 3.4. Human preference reflects a decrease in hallucinations with each additional round of self-augmentation, reaching saturation at Round-3. This observation aligns with the performance trends noted in Table 3 and GPT-4V and Gemini judges in Figure 2.
Model VSR random (%) $\uparrow$ VSR zero-shot (%) $\uparrow$
VILA0 73.5 63.1
VILA1 75.1 64.8
VILA2 76.8 65.8
VILA3 77.4 66.4
Table 7: Accuracies on the VSR benchmar.
Human Evaluation Win Rate (%) $\uparrow$
Later Rounds vs Original Caption 71.6
Later Rounds vs Round-1 54.6
Later Rounds vs Round-2 55.6
Later Round   vs Round-3 48.3
Table 8: Preferences for 200 images’ captions.
3.5 Improved Data Quality Matters More than Increased Computation
We next present additional ablations of VILA2’s self-augmentation loop in comparison to training additional epochs on the same data, with results shown in Table 9. We can observe that simply scaling up epochs with coarse image/text pairs yields no performance boosts, despite the extra training costs. In contrast, enhancing the quality of data, as demonstrated in VILA, provides a more rewarding path.
[…]
VILA0-Baseline 62.4 68.4 61.6 84.2 34.5 33.8
Train one extra epoch 62.5 68.7 61.9 84.0 34.4 33.9
VILA1 63.2 71.0 62.5 84.6 34.8 35.8
Train two extra epochs 62.3 69.0 61.7 83.9 34.4 33.7
VILA2 63.5 71.5 63.5 84.7 35.5 35.2
Table 9: Comparison between training additional epochs on the same data and training additional epochs with self-augmentation. Models do not benefit from more computations on identical data.
3.6 Efficiency and Effectiveness of VILA2
Cost Analysis – Labeling. Data quantity and quality are critical factors in model training. While VILA2 involves three rounds of recaptioning, it is still far more cost-efficient than traditional human re-labeling. For example, a standard re-labeling on Amazon Turk costs 36 USD per 1k images, while VILA2 costs only 0.12 USD per 1k images. The cost breakdown includes AWS pricing for H100 GPUs: USD 4.91 per hour for one H100, or USD 39.33 per hour for eight. With an inference speed of 10.6 images per second per H100, VILA2 processes around 38,340 images per hour, making it 300x cheaper and significantly faster.
Cost Analysis – Training. As shown in Table 10, VILA2 achieves better accuracy with significantly less data (51M) compared to other works like MM1 (>2B) and Idefics2 (>600M). Even with three rounds of iteration, VILA2 remains more cost-effective in terms of training computation. Further, VILA2 is a one-time cost that can be leveraged for training multiple models. Once recaptioning is complete, this data can be shared with the community, reducing the need for expensive data pipelines.
[…]
MM1-7B-Chat McKinzie et al. (2024) 7B 300M 720 $>$2B 82.8 72.6 72.8 72.3 64.0 81.5 42.1 35.6
Idefics2-8B Laurençon et al. (2024) 8B 400M 320 $>$600M 81.2 73.0 76.7 37.7
VILA2-8B (ours) 8B 400M 196 51M 82.9 87.6 73.4 76.6 66.1 86.6 50.0 38.3
Table 10: Comparison of multimodal methods across benchmarks, with different settings of large language model parameters (LLM), vision tower parameters (VT), number of tokens per image (# TPI), and pre-training data size (PT).
4 Related Work & Limitations
Visual language models (VLM). Visual language models have rapidly progressed in recent years  (Radford et al., 2021; Li et al., 2022a; Dai et al., 2023; Liu et al., 2023c; Ye et al., 2024; Cheng et al., 2024). The success mainly comes from pretraining visual and language models on the internet-scale data. Kosmos-2 (Peng et al., 2023) and PaLI-X (Chen et al., 2023c) largely scaled the pretraining data by pseudo-labeling bounding boxes from performant open-vocabulary object detectors (GLIP (Li et al., 2022b) and OWL-v2 (Minderer et al., 2024), respectively). They examined that strong perception capabilities such as object detection and OCR translate to better high-level reasoning tasks like visual question-answering (VQA).
Contributions & Novelty. Our work expands the horizon of data-scaling through our self-augmenting paradigm. ShareGPT4V (Chen et al., 2023b) applied a single round of recaptioning by distilling from GPT-4V. In contrast, we focus on a more general approach of using VLM to augment VLM itself without relying on commercial APIs or distilling from larger models. We provide 1) a detailed analysis of self-augmentation, covering prompt templates, iteration rounds, saturation points; 2) a practical method that uses specialist-augmentation to continually improve; 3) curated datasets that can be reused for future research. Our solution efficiently enhances SOTA VLM performance without requiring extra data or expensive APIs for closed-source models.